Suppose that the ages of undergraduate students at MIT is independently and normally distributed with a
mean of 19 and a standard deviation of 2. Suppose a first sample of four MIT students and a second sample
of four MIT students are selected at random.

1. What is the distribution of the sample mean, X, of the first four students? 2. What is the distribution of the sample mean, Y, of the other four students? 3. What is the distribution of 2X- Y?
Answer #1

1)as parent population is normal

Distribution of Xbar will be normal with mean \muxbar=19

and variance \sigma2xbar =\sigma2/n =(22/4)=1

or standard error of mean =\sigmaxbar =(1)1/2 =1

2)

as above:

Distribution of Ybar will be normal with mean \muybar=19

and variance \sigma2ybar =\sigma2/n =(22/4)=1

or standard error of mean =\sigmaybar =(1)1/2 =1

3)

distribution of 2Xbar-Ybar will be normal with

mean =2*19-19=19

and variance =(2)2 *\sigma2xbar+\sigma2ybar =4*1+1 =5

or std deviation =(5)1/2 =2.236